Inequality Problem:
A cell phone company offers a contract for which the cost (C), in dollars, of telephoning time is given by C = 0.25(t - 500)+ $48.95 where it is assumed that that the time(t) is greater than or equal to 500 minutes. What times in minutes will keep cost between $95.95 and $145.45?

Answer by stanbon(75887)

A cell phone company offers a contract for which the cost (C), in dollars, of telephoning time is given by C = 0.25(t - 500)+ $48.95 where it is assumed that that the time(t) is greater than or equal to 500 minutes. What times in minutes will keep cost between $95.95 and $145.45?
-----------------
95.95 <= 0.25(t - 500) + 48.95 <= 145.45
Subtract 48.95 along the line to get:
47 <= 0.25(t-500) <= 96.5
Multiply thru by 4 to get:
188 <= t-500 <= 386
Add 500 along the line to get:
688 <= t <= 886
=====================
Cheers,
Stan H.